OpenStudy (anonymous):

anyone good with algebra 2? i need help..

OpenStudy (accessdenied):

Do you have a particular question we can help you with? :)

OpenStudy (anonymous):

yea im having a hard time understanding quadratic equations

OpenStudy (accessdenied):

Anything more specific? A general quadratic equation is in the form, \(ax^2 + bx + c = 0\). What are you currently doing with these; factoring, solving, or completing squares? Or something more?

OpenStudy (anonymous):

uhm here lemme show you what im doing

OpenStudy (anonymous):

i just wanna know how to do the warm upsa

OpenStudy (accessdenied):

Are these supposed to be written that way? x^2 - 4x - 5 = 0x^2 - 4x - 5 = 0 8x^2 + 40x + 50 = 08x^2 + 40x + 50 = 0

OpenStudy (anonymous):

yea but they are 5 different questions

OpenStudy (accessdenied):

It is confusing that we have two equality signs per question, and some don't even equal the same thing. I can't decipher what the purpose of it...

OpenStudy (anonymous):

the first question is x^2-4x-5=0 and i dont know what to do

OpenStudy (accessdenied):

So the first step says to identify the a, b, and c values. That is coming from: \(ax^2 + bx + c = 0 \) \(\color{red}ax^2 + \color{green}bx + \color{blue}c = 0 \) \(\color{red}1x^2 \color{green}{- 4}x \color{blue}{- 5} = 0 \) Can you see how these compare so closely?

OpenStudy (accessdenied):

If there is no coefficient, we can assume it is 1. So for that x^2, it is equally expressed as 1x^2.

OpenStudy (accessdenied):

So for the first one, can you see from how these two equations compare that a=1, b=-4, and c=-5? I will give this one for a good reference point because there are others for you to try on your own as well.

OpenStudy (anonymous):

i see that, and is there a way to solve it or does it just stay 1x2−4x−5=0

OpenStudy (accessdenied):

We could technically solve for x if we needed (factoring looks applicable, also quadratic formula), but the questions do not ask us for the actual solutions! So no need to do the extra work!

OpenStudy (anonymous):

Awesome Sauce! thank you so much ^-^

OpenStudy (accessdenied):

We were asked to complete the following: (a) Identify a, b, and c. Which we did for the first one. (b) Find the discriminant, or in other words take the a, b, and c values from above and apply them to a formula: b^2 - 4ac. (c) Use that value from (b) to decide the number of solutions. The discriminant comes from the Quadratic Formula: \( x = \dfrac{ -b \pm \sqrt{ \color{goldenrod}{b^2 - 4ac}}}{2a} \leftarrow There\ it\ is\) The square root argument decides the number and type of solutions. If that discriminant is positive, we have -b + sqrt(#) and -b - sqrt(#). Two real solutions. If the discriminant is 0, the square root is 0; so we have -b + or - 0, or just -b. One real solution. If it is negative discriminant, we have to break out the complex numbers and we get two complex solutions!

OpenStudy (anonymous):

so complex is imaginary? so my answer for number one would be a- 1x2−4x−5=0 b- x= \[-b \pm \sqrt{b ^{2}}-4ac/2a\] c- two real solutions ?

OpenStudy (accessdenied):

Yeah, for the most part. Technically, complex is a better term than imaginary because complex means a real number added to an imaginary number, which is strictly something with i. But if they don't worry about that detail you shouldn't need to either. For (a) we specifically list our values of a, b, and c. Like, the equation is what they gave us, but we want to provide what is called a, b, or c? a=1, b=-4, or c=-5. So our answer would literally be just that: (a) a=value, b=value, and c=value. For (b) the values from (a) are used in the equation b^2 - 4ac. From above, we had a=1, b=-4, and c=-5. Plug those in: Our answer would be something like... Discriminant: b^2 - 4ac = (-4)^2 - 4*(1)*(-5) = simplify this to get the answer.

OpenStudy (anonymous):

so is the discriminant -4?

OpenStudy (accessdenied):

Not quite, the subtraction and minus sign combine to make a positive, right? :)

OpenStudy (anonymous):

Righto oops. i have been cramming all day because i have to do like 200 something classes before june 13th so im frettin. so it would be 26?

OpenStudy (accessdenied):

16 + 20 = 36, which I think you meant to put. That is correct.

OpenStudy (anonymous):

yea cx thank you so much this helped so much ^-^

OpenStudy (accessdenied):

Glad to help! And good luck on all those classes! That seems difficult to keep up with. o.o

OpenStudy (anonymous):

thank you haha ima need it. yea its mainly because i switched over cyber schools, and the one school i failed with a 64 in chem and a 63 in algebra 2 for the first semester, but in this school passing is a 62. and the school wouldnt give me passing for them so now i have to do both 1st and 2nd semesters for those two classes plus history, english, and some drawing classes.

OpenStudy (accessdenied):

Ugh, yea, very difficult indeed! Well, I hope you can get through it! I'll be here at least a bit every day if you wanted to tag me for a question (@AccessDenied), although sometimes I just leave the site open while I am gone. But I will try to help you when you need it! And I hope others will also do the same. :D

OpenStudy (anonymous):

thank you so much. im gonna go try the second one and come back to you and show you what i got and se if i got it right.

OpenStudy (anonymous):

for the second one we use the b^2-4ac and we plug in the stuff from a?

OpenStudy (accessdenied):

Yep, we will always take those values of a, b, and c for use in the formula. b^2 - 4ac.

OpenStudy (anonymous):

okay let me go through this and see what i get ^.^

OpenStudy (anonymous):

mehh i think i did somethink wrong Dx

OpenStudy (anonymous):

i keep getting this b^2-4ac=(-4)^2-4*(8)*(50)= -1592

OpenStudy (anonymous):

i dont know if -1592 is right or?

OpenStudy (accessdenied):

The original equation is: 8x^2 - 4x + 50 = 0 ?

OpenStudy (anonymous):

yeap

OpenStudy (accessdenied):

Not 8x^2 + 40x + 50 = 0 ? I thought I read that on the sheet earlier.

OpenStudy (anonymous):

Oh it is 40 xD my bad again long day lemme retry this

OpenStudy (anonymous):

now it comes out to 0 .-. is that right?

OpenStudy (accessdenied):

That looks better!

OpenStudy (anonymous):

so itd be one real solution correct?

OpenStudy (accessdenied):

That is correct.

OpenStudy (anonymous):

Yayyyy cx

OpenStudy (accessdenied):

I guess I should advise, in case you didn't already, to fix part (a) if you did put b=-4 instead of b=40. That would be silly to mix up there. :)

OpenStudy (anonymous):

there you go again helpin me x'D

OpenStudy (anonymous):

so for the second one b would be -223?

OpenStudy (anonymous):

third*

OpenStudy (accessdenied):

2x^2 + x + 28 = 0 a=2, b=1, and c=28 So 1^2 - 4*2*28 = -223. Looks good!

OpenStudy (anonymous):

okie dokie i think i got it from here, ill check in when im completly done.

OpenStudy (accessdenied):

Ohh I finally figured out what I was seeing that looked so weird. For some reason I am seeing the equation written twice on this document, side by side without spacing: \(\color{red}{x^2 -4x - 5 = 0} \color{blue}{x^2 - 4x - 5 = 0} \) \(x^2 - 4x - 5 = 0x^2 - 4x - 5 = 0 \) ??? That makes this seem a lot easier. lol.

OpenStudy (anonymous):

yea i think because when the teachers send it, its sent sideways idk why though cx

OpenStudy (anonymous):

how do we find out how many solutions it is again? Dx

OpenStudy (accessdenied):

Using the discriminant: b^2 - 4ac > 0 ==> 2 real solutions b^2 - 4ac = 0 ==> 1 real solution b^2 - 4ac < 0 ==> 2 (imaginary) solutions

OpenStudy (anonymous):

wow. that just made it 1000 times easier cx

OpenStudy (anonymous):

1) A|a=1 b=-4 c=-5 B|b^2 - 4ac = (-4)^2 - 4*(1)*(-5) = 36 C| two imaginary solutions. 2) A| a=8 b=40 c=50 B| b^2-4ac=(-40)^2-4*(8)*(50)=0 C| one real solution 3) A| a=2 b=1 c=28 B| b^2-4ac= (1)^2-4*(2)*(28)=-223 C|two real solutions. 4) A| a=6 b=-2 c=5 B| b^2-4ac=(-2)^2-4*(6)*(5)=-116 C|two real solutions. 5) A| a=1 b=8 c=-16 B| b^2-4ac=(8)^2-4*(1)*(-16)= -48 C|two real solutions. hows this?

OpenStudy (accessdenied):

First thing I notice is you flipped around the condition for two real and two imaginary solutions. The discriminant (b^2 - 4ac) > 0 means two REAL solutions, and < 0 means two IMAGINARY solutions. For #1, you had discriminant of 36, but that means two REAL solutions because it is greater than 0.

OpenStudy (anonymous):

so is everything else correct then?

OpenStudy (accessdenied):

The other thing is specifically for #5. Notice the equation is written: \(x^2 = 8x - 16 \) That is actually not in the form \(ax^2 + bx + c = 0\) yet! We need to get 0 on one side of the equation first, by moving everything else over with x^2. We subtract 8x and add 16 to both sides. \(x^2 \color{green}{-8x + 16} = 0 \)

OpenStudy (anonymous):

Oh ok ok so ima put that and actually use that to solve ok cx

OpenStudy (accessdenied):

Yep. That is a really common mistake; they always try to pull that on students, mixing up the equation. You just need to check and make sure there is one side that is 0. That's not too much to inspect. :)

OpenStudy (anonymous):

5) = x^2-8x+16=0 A| a=1 b=-8 c=16 B| b^2-4ac= (-8)^2-4*(1)*(16)= 0 C| one real solution this is what i got

OpenStudy (accessdenied):

Looks good now!

OpenStudy (anonymous):

thank you so much ^-^

OpenStudy (accessdenied):

Glad to help!

OpenStudy (anonymous):

@AccessDenied are you on Dx